Sweeney originally contributed $302,000 in cash for a one-fourth interest in the Gilbert LLC. During the several years that Sweeney was a member of the LLC, his share of the LLC's income was $181,200 and he withdrew $120,800 cash. The LLC's liabilities are $151,000, of which Sweeney's share is $37,750. The LLC has $75,500 of hot assets, of which Sweeney's share is $18,875. Sweeney sells his LLC interest to Jana for $398,640 cash, with Jana also assuming Sweeney's share of the LLC's liabilities. How much is Sweeney's gain on the sale, and what is its character?
Answer #1

398640+37750 = 436390

302000+181200-120800 = 362400

362400+37750 = 400150

gain on sale = 400150-398640 = 1510